OpenStudy (anonymous):

if sin theta=3/5 and theta is in quad 2 the exact form of sin (theta+ pi/6) is ..... I had already asked this question but still don't understand it

OpenStudy (butterflydreamer):

\[\sin (\theta + \alpha) = \sin \theta \cos \alpha + \sin \alpha \cos \theta\]

OpenStudy (butterflydreamer):

therefore: \[\sin (\theta + \frac{ \pi }{ 6}) = \sin \theta \cos \frac{ \pi }{ 6 } + \sin \frac{ \pi }{ 6 } \cos \theta\]

OpenStudy (butterflydreamer):

and then you want to use your exact value table to help you out here since they want the final answer in exact form (this means including radicals) :)

OpenStudy (butterflydreamer):

 also note: \[\pi/6 = 180/6 = 30\]